# Probabilistic Counters for Privacy Preserving Data Aggregation11footnotemark: 1

### v5

@@ 3 Differential Privacy Preliminaries
In this section, we briefly recall differential privacy. For more details see, e.g. [5]. We denote the set of (positive) natural numbers by $\mathbb{N}$ and the set of all integers by $\mathbb{Z}$. Moreover, let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $a,b\in\mathbb{Z}$ let us define a discrete interval $[a,b]\cap\mathbb{Z}$ by $[a:b]$. We also define $[n]=\{1,2,\ldots,n\}$ for $n\in\mathbb{N}$. We assume that there exists a trusted curator who holds, or securely obtains, the data of individuals in a (possibly distributed) database $x$. Every row of $x$ consists of the data of some individual. By $\mathcal{X}$, we denote the space of all possible rows. The goal is to protect the data of every single individual, even if all users except one collude with an adversary to breach the privacy of this single, uncorrupted user. On the other hand, the curator is responsible for producing a release – a possibly accurate response to a requested query. This response is then released to the public, which is allowed to perform statistical analysis on it. The differential privacy is, by design, resilient to post-processing attacks, so even if the adversary obtains the public release, he will not be able to infer anything about specific individuals participating in this release.
For simplicity, we interpret databases as their histograms in $\mathbb{N}_{0}^{|\mathcal{X}|}$, so we may just focus on unique rows and the numbers of their occurrences.
@@
where $x_{i}$ and $y_{i}$ denote the numbers of occurrences of an item (an individual) $i$ in the databases $x$ and $y$ respectively.
@@
A privacy mechanism is a randomized algorithm used by the curator that takes a database as input and produces the output (the release) using randomization.
@@
A randomized algorithm $\mathcal{M}$ with domain $\mathbb{N}^{|\mathcal{X}|}$ is ($\varepsilon,\delta$)-differentially private (or ($\varepsilon,\delta$)-DP), if for all $\mathcal{S}\subseteq$ $\mathcal{M}$) and for all $x,y\in\mathbb{N}^{|\mathcal{X}|}$ such that $\left\|x-y\right\|_{1}\leqslant 1$ the following condition is satisfied:
@@
(Laplace noise) In the central model, a standard, widely used mechanism with $(\varepsilon)$-DP property is the so-called Laplace noise. A variable $X$ has Laplace distribution with parameter $\lambda$ (denoted as $X\sim\mathcal{L}(\lambda)$), if its probability density function is
@@
Let $c(x)$ be the number of rows in $x$, that satisfy a given property. Note that $c$ in the differential privacy literature is usually referred to as count query. Imagine that an aggregating mechanism is defined as follows: $\mathcal{M}(x)=c(x)+\mathcal{L}(\varepsilon^{-1})$. Then $\mathcal{M}$ is $(\varepsilon)$-DP (for more precise properties of Laplace noise, see [5]).
@@ 4 Probabilistic Counters — preliminaries
This paper focuses on probabilistic counters, denoted further by $M$. The notion of the probabilistic counter is ambiguous in literature. It is a stochastic process that can be interpreted as a mechanism defined on the space of all possible inputs, which should estimate some goal value in some sense. The exact definition of this approximation is not crucial from DP-point of view, thence we do not consider it in this paper.
Each increase of the data source being counted by the probabilistic counter is called an increment request. Due to the randomized nature of probabilistic counters, each may change the value of the counter, but not necessarily. We will further indicate the single increment request by ${}^{\prime}1^{\prime}$. For the sake of generality, we also assume that the counter can get as an input ${}^{\prime}0^{\prime}$, and in such case, it simply does nothing. This is useful for real-life scenarios, e.g., data aggregation (see Section 5). Obviously, only increment requests impact the counter’s final result; hence, we indicate the counter’s value after $n$ increment requests by $M_{n}$, and we are not considering the number of the rest of the rows.
In Figure 1, one can see a graphical representation of the probabilistic counter. As mentioned, increment requests are indicated by ${}^{\prime}1^{\prime}$ and other rows by ${}^{\prime}0^{\prime}$ input. The dice depicts the randomness. The $\mathbf{X}$-mark stands for no action.
@@
We emphasize that the probabilistic counter depends on the number of increment requests. We want to show that if we reveal its final value, then it does not expose any sensitive data about any single record. Moreover, note that if $x$ and $y$ differ only by one ${}^{\prime}0^{\prime}$ input, then $\mathbb{P}\left(M(x)\in\mathcal{S}\right)=\mathbb{P}\left(M_{n}\in\mathcal{S}\right)=\mathbb{P}\left(M(y)\in\mathcal{S}\right)$, where $n$ is the number of increment requests both for $x$ and $y$. See that then the condition in Definition 2 is trivially fulfilled. Hence for our convenience, in this paper, we use to mark only the number $n$ of increment requests provided by individuals when talking about the probabilistic counter $M_{n}$.
@@
Note that, for our setting, Fact 1 is fully compatible with the intuition of regular differential privacy (Definition 2). Indeed, Fact 1 can be easily derived from the observation that any set $B\subset A$ is a disjoint union of $B\cap S_{n}$ and $B\cap S_{n}^{\prime}$.
Remark that $\varepsilon$ and $\delta$ in Fact 1 can be also treated as functions of a parameter $n$, i.e. with respect to the number of increment requests. Thence, we can consider either differential privacy of $M_{n}$ variable, when $n$ is known or ($\varepsilon(n),\delta(n))$-DP of probabilistic counter $M$. The second variant let us provide a precise dependence of privacy parameters of the counter as the number of increment requests gets large.
@@ 4.1 Morris Counter
We begin with a short description of Morris Counter (originally referred to as an approximate counter [1, 2]). Let us fix $a>1$. Algorithm 1 is a very simple pseudocode of Morris Counter [1].
$M\leftarrow 1$;
@@ 5             $M\leftarrow M+1$;
Morris Counter is a Markov process $(M_{n},n\in\mathbb{N}_{0})$ which satisfies:
@@
Note that Definition 3 can be derived directly from a run of Algorithm 1. From now on let $\mathbb{P}\left(M_{n}=l\right)=p_{n,l}$. Directly from Definition 3 we get the following recursion:
@@
The accuracy of Morris Counter has been thoroughly analyzed in various classical papers. The first detailed analysis was proposed by Ph. Flajolet in [2]. In this part we present the essence of theorems presented in this paper, which will be useful later on.
@@
which will emerge as a crucial point of our further considerations of this Markov process in terms of differential privacy in this section. Let us mention, that usually $[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4]\subseteq[n+1]$, so one may think that $I_{n}$ are symmetric discrete intervals of length $8$, centered at $\lceil\log(n)\rceil$.
The lion’s share of applications of Morris Counter is based on counting a number of occurrences, that is, the number of increment requests. In order to estimate this value, we may use  (3) and simply obtain $\mathbb{E}\left(2^{M_{n+1}}\right)=\mathbb{E}\left(2^{M_{n}}\right)+1$, so together with the assumption $M_{0}=1$ we obtain
@@
Hence $22^{M_{n}}-$ 2 is an unbiased estimator of the number of increments $n$. Remark that $n$ can be saved in $\lceil\log(n)\rceil$ bits. On the other hand, Fact 2 shows that on average, $\log(\log(n))+O(1)$ bits are required to store $M_{n}$. As announced earlier, this is the crucial advantage of Morris Counter. Moreover, analogously to (5) we may obtain
@@
Formulas (5) and (6) will be used in the analysis of data aggregation example in Section 5.
@@ 4.2 MaxGeo Counter
We begin with a short description of MaxGeo Counter. Algorithm 2 shows its pseudocode. Informally, for each increment request, the server has to generate a random variable from the geometric distribution $\mathrm{Geo}({1}/{2})$ (ranged in $\mathbb{N}$). The final result is the maximum taken over all these generated random variables.
$C\leftarrow 1$;
@@ 4       $C\leftarrow\max\{C,r\}$;
The expectation and variance of the maximum of $n$ i.i.d. geometric variables have already been analyzed in the literature. For instance, Szpankowski and Rego [8] provided exact formulas for such variables’ expected value and variance. However, they are impractical for applications for large $n$. Hence they also provided asymptotics (here, for a maximum of $n$ independent $\mathrm{Geo}(1/2)$ distributions): $\mathbb{E}\left(M_{n}\right)=\log(n)+O(1)$ and $\mathrm{Var}(M_{n})=\log(n)+O(1)$ and thus, similarly to Morris Counter, there are only $\log(\log(n))+O(1)$ bits required on average to save MaxGeo Counter after $n$ increment requests.
@@ 4.3 General Probabilistic Counting with Stochastic Averaging
Here we recall briefly a General Probabilistic Counting with Stochastic Averaging algorithm, based on the original idea from [25]. Assume that there are $m$, initially empty lots related to some independent copies of some probabilistic counter. For each increment request, we connect it with one of the groups uniformly at random. Finally, we perform incrementation requests separately and independently for each lot, obtaining $M[1],M[2],\ldots,M[m]$.
Without delving into details, for original PCSA algorithm, $\mathbb{E}\left(M_{n}\right)\approx\log(\varphi n)$, where $M_{n}$ is a value of a specific probabilistic counter connected with PCSA after $n$ increment requests and $\varphi$ is some magic constant. If we denote the mean of these $m$ counters after the total number of increment requests $n$ by $\sigma_{n}(m)$, then we may introduce the statistic:
@@
An important conclusion is that we may apply the idea of original PCSA in general, for any probabilistic counters.
@@ 4.4 LogLog counter
In [44] a LogLog algorithm was proposed. It is based on $m=2^{k}$ counters $(M[j])_{j=1}^{m}$, where $k>0$. We may interpret this algorithm in the context of probabilistic counters. In such a scenario, it takes a hashed value (binary sequence) as an input on every increment request. The first $k$ bits of the hash determine which of the $m$ counters should be incremented (the index $j$ is chosen as the decimal representation of the sequence restricted to these first $k$ bits; hence it translates the increment request of the LogLog counter to the increment request of one of the $m$ internal counters). Consider the first non-zero bit of a tail of the sequence (starting from $(k+1)$-th bit). Its position $R$ in this tail follows the $\mathrm{Geo}\left(\frac{1}{2}\right)$ distribution assuming the uniform distribution of the input sequences. If $R>M[j]$, then $M[j]$ should become $R$. Otherwise it does not change.
Therefore, LogLog counter is, in fact, a general PCSA that uses $m$ MaxGeo counters. It can be used to estimate the cardinality of the increment requests $n$ using the following estimator:
@@
We are interested in the expectation and accuracy of the cardinality estimator, which can be briefly described as follows: $\mathbb{E}\left(\mathrm{LogLog}_{n}^{(m)}\right)\approx n$ and $\mathrm{Var}(\mathrm{LogLog}_{n}^{(m)})\approx\frac{1.69n^{2}}{m}$ (check [44] for more details). However, in order to control the LogLog counter, on average about $m\log\log\left(\frac{n}{m}\right)$ bits of memory are needed.
An interesting fact is that the expectation of single MaxGeo counter is logarithmic and the estimator of $n$ is of the form $Cm2^{\sigma_{n}(m)}$, for some constant $C$ (just like in the case of original PCSA).
@@ 4.5 HyperLogLog
The maximum of geometric variables is also used as a primitive in well known HyperLogLog algorithm (see [9]). Therefore its privacy properties are important both from the theoretical and practical point of view. Essentially, in HyperLogLog we perform the general PCSA algorithm, but the final estimation is somehow different:
@@
where $\alpha_{k}$ is a constant dependent only on $k$ (see [9] for more details). It is worth noting that HyperLogLog related algorithms (mentioned in Section 1) are the best-known procedures designated for cardinality estimation, and they are close to optimum [45]. According to [9], for $m=2^{k}$, where $k\geqslant 4$,
@@ 5 Privacy-Preserving Survey via Probabilistic Counters
In this section, we present an example scenario for data aggregation using probabilistic counters. We assume there is a server (alternatively, we call it aggregator) and a collection of nodes (e.g., mobile phone users), and we want to perform a boolean survey with a sensitive question. Namely, each user sends ${}^{\prime}0^{\prime}$ if his/her answer is no and ${}^{\prime}1^{\prime}$ if the answer is yes. We assume that the connections between users and the server are perfectly secure and the data can safely get to the trusted server. This can be performed using standard cryptographic solutions. The goal of the server is to publish the sum of all ${}^{\prime}1^{\prime}$ responses in a privacy-preserving way. Such a goal could obviously be achieved by simply collecting all the data and adding an appropriately calibrated, Laplace noise (see [5]). However, we aim to show that probabilistic counters have inherently sufficient randomness to be differentially private without any auxiliary randomizing mechanism.
@@
    when all data is processed, the value of the counter is released to the public.
Note that we assume that the Adversary has access only to the released value. We also released just the counter value itself, which does not estimate ${}^{\prime}1^{\prime}$ responses. Such estimation is a function of released value, which is different for Morris or MaxGeo Counter. There can also be various ways to estimate the actual number using counter value. However, this does not matter for our case, as differential privacy is, conveniently, fully resilient to post-processing (see [5]). The graphical depiction of our considered scenario is presented in Figure 2.
@@
Our assumptions about the Adversary are the same as in most differential privacy papers. Namely, he may collude with any subset of the participants (e.g., all except the single user whose privacy he wants to breach). On the other hand, the aggregator is trusted. See that even though we have a distributed system in mind, this is, in fact, a central differential privacy scenario. We do not assume pan-privacy. It means that the algorithm’s internal state is not subject to the constraints of differential privacy. Obviously, if the Adversary would know the internal state of the counter at any time or could observe whether, after receiving data from a specific user, the server has to perform computations to potentially increment the counter (implying a ${}^{\prime}1^{\prime}$ response) or not, he would easily violate the privacy. We also do not assume privacy under continual observation. The survey is not published iteratively but one time only after it is finished. To sum it up, the Adversary cannot
@@
    collude with any subset $C$ of the participants (e.g. know their data or make them all send ${}^{\prime}0^{\prime}$ to the server (intentially)) in order to breach the privacy of user not belonging to $C$,

### v6

@@ 3 Differential Privacy Preliminaries
In this section, we briefly recall differential privacy. For more details, see, e.g. [18]. We denote the set of (positive) natural numbers by $\mathbb{N}$ and the set of all integers by $\mathbb{Z}$. Moreover, let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For $a,b\in\mathbb{Z}$ let us define a discrete interval $a,b]\cap\mathbb{Z}$ by $[a:b]$. We also define $n]=\{1,2,\ldots,n\}$ for $n\in\mathbb{N}$. We assume that there exists a trusted curator who holds, or securely obtains, the data of individuals in a (possibly distributed) database $x$. Every row of $x$ consists of the data of some individual. By $\mathcal{X}$, we denote the space of all possible rows. The goal is to protect the data of every single individual, even if all users except one collude with an adversary to breach the privacy of this single, uncorrupted user. On the other hand, the curator is responsible for producing a release – a possibly accurate response to a requested query. This response is then released to the public, who is allowed to perform a statistical analysis on it. The differential privacy is, by design, resilient to post-processing attacks, so even if the adversary obtains the public release, he will not be able to infer anything about specific individuals participating in this release.
For simplicity, we interpret databases as their histograms in $\mathbb{N}_{0}^{|\mathcal{X}|}$, so we can focus only on unique rows and the numbers of their occurrences.
@@
where $x_{i}$ and $y_{i}$ denote the numbers of occurrences of an item (an individual) $i$ in the databases $x$ and $y$, respectively.
@@
A privacy mechanism is a randomised algorithm used by the curator that takes a database as input and produces the output (the release) using randomisation.
@@
A randomized algorithm $\mathcal{M}$ with domain $\mathbb{N}^{|\mathcal{X}|}$ is ($\varepsilon,\delta$)-differentially private (or ($\varepsilon,\delta$)-DP), if for all $\mathcal{S}\subseteq$ Range($\mathcal{M}$) and for all $x,y\in\mathbb{N}^{|\mathcal{X}|}$ such that $\left\|x-y\right\|_{1}\leqslant 1$ the following condition is satisfied:
@@
(Laplace noise) In the central model, a standard and widely used mechanism with the $(\varepsilon)$-DP property is the so-called Laplace noise. A variable $X$ has Laplace distribution with parameter $\lambda$ (denoted as $X\sim\mathcal{L}(\lambda)$), if its probability density function is
@@
Let $c(x)$ be the number of rows in $x$, which satisfy a given property. Note that $c$ in the differential privacy literature is usually referred to as count query. Imagine that an aggregating mechanism is defined as follows: $\mathcal{M}(x)=c(x)+\mathcal{L}(\varepsilon^{-1})$. Then $\mathcal{M}$ is $(\varepsilon)$-DP (for more precise properties of Laplace noise, see [18]).
In the privacy analysis of large-scale distributed protocols, two types of approaches are typically distinguished: event/record-level (e.g., [51, 42]), where information about a specific event/record is protected, and user-level privacy (e.g., [34]), which protects the privacy of the users themselves. Generally, the latter is a stronger model (see the discussion in [34]).
Fundamentally, the model analysed in our work pertains to event-level privacy. That is, we protect information regarding whether a counter incrementation event occurred. This corresponds to a scenario in which a participant can perform a certain action only once (for example, liking a specific message on typical social media platforms). These results can be directly adapted to scenarios where a single user may be associated with multiple counter increments using standard methods. Naturally, the strength of the privacy guarantee in such a framework must depend heavily on additional assumptions (such as constraints on the number of events or the associations between users).
@@ 4 Probabilistic Counters — preliminaries
This paper focuses on probabilistic counters, further denoted by $M$. The notion of a probabilistic counter is ambiguous in the literature. It is a stochastic process that can be interpreted as a mechanism defined on the space of all possible inputs, which should estimate some goal value in some sense. The exact definition of this approximation is not crucial from the DP-point of view; thus we do not consider it in this paper.
Each increase in the data source counted by the probabilistic counter is called an increment request. Due to the randomised nature of probabilistic counters, each may change the value of the counter, but not necessarily. We will also indicate the single increment request by $^{\prime}1^{\prime}$. For the sake of generality, we also assume that the counter can get as an input $^{\prime}0^{\prime}$, and in such a case it simply does nothing. This is useful for real-life scenarios, e.g., data aggregation (see Section 5). Obviously, only increment requests impact the counter’s final result; hence, we indicate the counter’s value after $n$ increment requests by $M_{n}$, and we are not considering the number of the rest of the rows.
In Figure 1, one can see a graphical representation of the probabilistic counter. As mentioned, increment requests are indicated by $^{\prime}1^{\prime}$ and other rows by $^{\prime}0^{\prime}$ input. The dice represent the randomness. The $\mathbf{X}$-mark indicates that there is no action.
@@
We emphasise that the probabilistic counter depends on the number of increment requests. We want to show that if we reveal its final value, then it does not expose any sensitive data about any single record. Moreover, note that if $x$ and $y$ differ only by one input $^{\prime}0^{\prime}$, then $\mathbb{P}\left(M(x)\in\mathcal{S}\right)=\mathbb{P}\left(M_{n}\in\mathcal{S}\right)=\mathbb{P}\left(M(y)\in\mathcal{S}\right)$, where $n$ is the number of increment requests for both $x$ and $y$. See that then the condition in Definition 2 is trivially fulfilled. Hence, for our convenience, in this paper, we use the marking of only the number $n$ of increment requests provided by individuals when talking about the probabilistic counter $M_{n}$.
@@
Note that, for our setting, Fact 1 is fully compatible with the intuition of regular differential privacy (Definition 2). Indeed, Fact 1 can easily be derived from the observation that any set $B\subset A$ is a disjoint union of $B\cap S_{n}$ and $B\cap S_{n}^{\prime}$.
Remark that $\varepsilon$ and $\delta$ in Fact 1 can also be treated as functions of a parameter $n$, i.e. with respect to the number of increment requests. Thence, we can consider the differential privacy of the variable $M_{n}$, when $n$ is known, or ($\varepsilon(n),\delta(n))$-DP of the probabilistic counter $M$. The second variant lets us provide a precise dependence of privacy parameters of the counter as the number of increment requests gets large.
@@ 4.1 Morris Counter
We begin with a short description of the Morris Counter (originally referred to as an approximate counter [36, 24]). Fix $a>1$. The algorithm 1 is a very simple pseudocode of the Morris Counter [36].
$1M\leftarrow$ 1;
@@ 5    $M\leftarrow M+1$;
The Morris Counter is a Markov process $(M_{n},n\in\mathbb{N}_{0})$ that satisfies:
@@
Note that Definition 3 can be derived directly from a run of Algorithm 1. From now on, let $\mathbb{P}\left(M_{n}=l\right)=p_{n,l}$. Directly from Definition 3 we get the following recursion:
@@
The accuracy of Morris Counter has been thoroughly analysed in various classical papers. The first detailed analysis was proposed by Ph. Flajolet in [24]. In this part, we present the essence of theorems presented in this paper, which will be useful later on.
@@
which will emerge as a crucial point of our further considerations of this Markov process in terms of differential privacy in this section. Let us mention that usually $\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4]\subseteq[n+1]$, so one may think that $I_{n}$ are symmetric discrete intervals of length $8$, centred at $\lceil\log(n)\rceil$.
The lion’s share of applications of Morris Counter is based on counting a number of occurrences, that is, the number of increment requests. In order to estimate this value, we may use  (3) and simply obtain $\mathbb{E}\left(2^{M_{n+1}}\right)=\mathbb{E}\left(2^{M_{n}}\right)+1$, so together with the assumption $M_{0}=1$ we obtain the following.
@@
Hence $2^{M_{n}}-2$ is an unbiased estimator of the number of increments $n$. Remark that $n$ can be saved in $\lceil\log(n)\rceil$ bits. On the other hand, Fact 2 shows that on average, $\log(\log(n))+O(1)$ bits are required to store $M_{n}$. As announced earlier, this is the crucial advantage of Morris Counter. Moreover, analogously to (5) we may obtain
@@
Formulas (5) and (6) will be used in the example of data aggregation analysis in Section 5.
@@ 4.2 MaxGeo Counter
We begin with a short description of MaxGeo Counter. Algorithm 2 shows its pseudocode. Informally, for each increment request, the server has to generate a random variable from the geometric distribution $\mathrm{Geo}({1}/{2})$ (ranged in $\mathbb{N}$). The final result is the maximum taken over all these random variables generated.
margin:
$1C\leftarrow$ 1;
@@ 4  $C\leftarrow\max\{C,r\}$;
The expectation and variance of the maximum of $n$ i.i.d. geometric variables have already been analysed in the literature. For instance, Szpankowski and Rego [47] provided exact formulas for the’ expected value and variance of such variables. However, they are impractical for large applications $n$. Hence, they also provided asymptotics (here, for a maximum of $n$ independent $\mathrm{Geo}(1/2)$ distributions): $\mathbb{E}\left(M_{n}\right)=\log(n)+O(1)$ and $\mathrm{Var}(M_{n})=\log(n)+O(1)$ and thus, similarly to the Morris Counter, there are only $\log(\log(n))+O(1)$ bits required on average to save the MaxGeo Counter after $n$ increment requests.
@@ 4.3 General Probabilistic Counting with Stochastic Averaging
Here we recall briefly a General Probabilistic Counting with Stochastic Averaging algorithm, based on the original idea from [23]. Assume that there are $m$, initially empty lots related to some independent copies of some probabilistic counter. For each increment request, we connect it to one of the groups uniformly at random. Finally, we perform incrementation requests separately and independently for each lot, obtaining the following. $M[1],M[2],\ldots,M[m]$.
Without delving into details, for the original PCSA algorithm, $\mathbb{E}\left(M_{n}\right)\approx\log(\varphi n)$, where $M_{n}$ is a value of a specific probabilistic counter connected with PCSA after $n$ increment requests and $\varphi$ is some magic constant. If we denote the mean of these counters $m$ after the total number of increment requests $n$ by $\sigma_{n}(m)$, then we may introduce the statistic:
@@
An important conclusion is that we may apply the idea of original PCSA in general to any probabilistic counters.
@@ 4.4 LogLog counter
In [13] a LogLog algorithm was proposed. It is based on $m=2^{k}$ counters $(M[j])_{j=1}^{m}$, where $k>0$. We may interpret this algorithm in the context of probabilistic counters. In such a scenario, it takes a hashed value (binary sequence) as input on every increment request. The first $k$ bits of the hash determine which of the $m$ counters should be incremented (the index $j$ is chosen as the decimal representation of the sequence restricted to these first $k$ bits; hence it translates the increment request of the LogLog counter to the increment request of one of the $m$ internal counters). Consider the first non-zero bit of a tail of the sequence (starting from $(k+1)$-th bit). Its position $R$ in this tail follows the $\mathrm{Geo}\left(\frac{1}{2}\right)$ distribution assuming the uniform distribution of the input sequences. If $R>M[j]$, then $M[j]$ should become $R$. Otherwise, it does not change.
Therefore, LogLog counter is, in fact, a general PCSA that uses $m$ MaxGeo counters. It can be used to estimate the cardinality of increment requests $n$ using the following estimator:
@@
We are interested in the expectation and accuracy of the cardinality estimator, which can be briefly described as follows: $\mathbb{E}\left(\mathrm{LogLog}_{n}^{(m)}\right)\approx n$ and $\mathrm{Var}(\mathrm{LogLog}_{n}^{(m)})\approx\frac{1.69n^{2}}{m}$ (see [13] for more details). However, in order to control the LogLog counter, on average about $m\log\log\left(\frac{n}{m}\right)$ bits of memory are needed.
An interesting fact is that the expectation of a single MaxGeo counter is logarithmic and the estimator of $n$ is of the form $Cm2^{\sigma_{n}(m)}$, for some constant $C$ (just as in the case of the original PCSA).
@@ 4.5 HyperLogLog
The maximum of geometric variables is also used as a primitive in the well known HyperLogLog algorithm (see [22]). Therefore, its privacy properties are important both from the theoretical and practical point of view. Essentially, in HyperLogLog we perform the general PCSA algorithm, but the final estimation is somehow different:
@@
where $\alpha_{k}$ is a constant dependent only on $k$ (see [22] for more details). It should be noted that HyperLogLog related algorithms (mentioned in Section 1) are the best-known procedures designated for cardinality estimation, and are close to optimum [30]. According to [22], for $m=2^{k}$, where $k\geqslant 4$,
@@ 5 Privacy-Preserving Survey via Probabilistic Counters
In this section, we present an example scenario for data aggregation using probabilistic counters. We assume that there is a server (alternatively, we call it aggregator) and a collection of nodes (e.g., mobile phone users), and we want to perform a boolean survey with a sensitive question. That is, each user sends $^{\prime}0^{\prime}$ if his answer is no and $^{\prime}1^{\prime}$ if the answer is yes. We assume that the connections between users and the server are perfectly secure, and the data can safely get to the trusted server. This can be performed using standard cryptographic solutions. The server’s goal is to publish the sum of all responses $^{\prime}1^{\prime}$ in a way that preserves privacy. This goal could obviously be achieved by simply collecting all the data and adding an appropriately calibrated Laplace noise (see [18]). However, we aim to show that probabilistic counters have inherently sufficient randomness to be differentially private without any auxiliary randomising mechanism.
@@
    when all data are processed, the value of the counter is released to the public.
Note that we assume that the Adversary has access only to the released value. We also released only the counter value itself, which does not estimate the responses of $^{\prime}1^{\prime}$. This estimation is a function of the released value, which is different for Morris or MaxGeo Counter. There can also be various ways to estimate the actual number using a counter value. However, this does not matter for our case, as differential privacy is conveniently fully resilient to post-processing (see [18]). The graphical representation of our scenario considered is presented in Figure 2.
@@
Our assumptions about the Adversary are the same as in most differential privacy papers. Namely, he may collude with any subset of the participants (e.g., all except the single user whose privacy he wants to breach). On the other hand, the aggregator is trusted. See that even though we have a distributed system in mind, this is, in fact, a central differential privacy scenario. We do not assume pan-privacy. This means that the algorithm’s internal state is not subject to the constraints of differential privacy. Obviously, if the Adversary knew the internal state of the counter at any time or could observe whether, after receiving data from a specific user, the server had to perform computations to potentially increment the counter (implying a $^{\prime}1^{\prime}$ response) or not, he would easily violate the privacy. We also do not assume privacy under continual observation. The survey is not published iteratively but once only after it is finished. In short, the Adversary cannot
@@
    collude with any subset $C$ of the participants (e.g., know their data or send them all $^{\prime}0^{\prime}$ to the server) in order to breach the privacy of the user not belonging to $C$,